OpenStudy (anonymous):

Write an equation of the line that passes through the points (1, 1) and (7, 4). A) y = 2x - 10 B) y = 2x - 1 C) y = (1/2)x + 5 D) y = (1/2)x + 1/2 Please help I am so confused

OpenStudy (solomonzelman):

First, find the slope of the line. \(\LARGE \color{black}{ \displaystyle {\rm m}=\frac{\color{blue}{{\rm y}_1}-\color{red}{{\rm y}_2}}{\color{green}{{\rm x}_1}-\color{darkgoldenrod}{{\rm x}_2}} }\) where \(\LARGE \color{black}{ \displaystyle {\rm m} }\) is the slope \(\Large\color{black}{ \displaystyle (\color{green}{{\rm x}_1}~,~~\color{blue}{{\rm y}_1}) }\) and \(\Large\color{black}{ \displaystyle (\color{darkgoldenrod}{{\rm x}_2}~,~~\color{red}{{\rm y}_2}) }\) are you points.

OpenStudy (solomonzelman):

in your case \(\LARGE \color{black}{ \displaystyle {\rm m}=\frac{\color{blue}{1}-\color{red}{4}}{\color{green}{1}-\color{darkgoldenrod}{7}} }\)

OpenStudy (solomonzelman):

what would your slope (m) be?

OpenStudy (anonymous):

3/6 or 1/2

OpenStudy (solomonzelman):

yes, that is correct, m=1/2

OpenStudy (solomonzelman):

Now you can use the point slope form to find the equation of the line. \(\large\color{black}{ \displaystyle {\rm y}-\color{purple}{{\rm y}_1} ={\rm m}\left({\rm x-\color{purple}{{\rm x}_1}}\right) }\)

OpenStudy (solomonzelman):

you can use any of the points, lets choose (1,1) \(\large\color{black}{ \displaystyle {\rm y}-\color{purple}{1} =\frac{1}{2}\left({\rm x-\color{purple}{1}}\right) }\)

OpenStudy (solomonzelman):

now, all you need is to simplify this expression.

OpenStudy (anonymous):

I am a little confused still sorry

OpenStudy (solomonzelman):

what exactly confuses u?

OpenStudy (solomonzelman):

well, as it is now it is also fine in a point slope form. However to make it one of your options you need a y-intercept form.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle {\rm y}-\color{purple}{1} =\frac{1}{2}\left({\rm x-\color{purple}{1}}\right) }\) \(\large\color{black}{ \displaystyle {\rm y}-\color{purple}{1} =\frac{1}{2}{\rm x}-\frac{1}{2}}\)

OpenStudy (solomonzelman):

then, add 1 to both sides, and you get?

OpenStudy (anonymous):

I understand the y- 1 + 1/2x = 1/2 but I am confused when I add 1 to each side

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle {\rm y}-1 =\frac{1}{2}{\rm x}-\frac{1}{2}}\) \(\large\color{black}{ \displaystyle {\rm y}-1\color{red}{+1} =\frac{1}{2}{\rm x}-\frac{1}{2}\color{red}{+1} }\) \(\large\color{black}{ \displaystyle {\rm y} =\frac{1}{2}{\rm x}+\frac{1}{2}}\)

OpenStudy (anonymous):

Now it makes better sense. I am sorry Thank you for the break down to help me understand it better